# RG theory of spontaneous stochasticity for Sabra model of turbulence

### v1

@@ VII Conclusion
The RG theory is formulated for discrete-time shell models. Such models are developed as approximations of the continuous-time Sabra model, converging to the latter as the approximation parameter $\varepsilon$ tends to zero. These discrete models are built on a dynamic space-time lattice and, from a numerical point of view, represent an adaptive finite-difference scheme. The key point is that these approximations exactly preserve scale symmetries and energy balance.
The RG theory applies to class of models, in which the regularization (dissipation and noise) terms are scale invariant. The ideal limit corresponds to $N\to\infty$, where $N$ is a regularization (cutoff) parameter. For each $N$, the stochastic evolution is described by a flow $\Phi^{(N)}$, which is a family of transition probabilities (Markov kernels) for all initial and boundary conditions. We derive the RG operator $\mathcal{R}$, which expresses the flow $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. This operator defines a dynamical system in the space of flows and associates the ideal limit with an attractor of the RG dynamics. Here the dynamics (operator $\mathcal{R}$) depends on nonlinear couplings of the ideal system, but does not depend on the regularization (dissipation and noise). A specific regularization is introduced into the RG dynamics through the initial condition, i.e. through the initial flow $\Phi^{(N_{0})}$.
@@
The RG approach to Eulerian spontaneous stochasticity was earler developed for discrete models on a self-similar (fractal) lattice [33, 31]. Our current work demonstrates that the same approach works for systems with continuous time such as the Sabra shell model. This is achieved by using a dynamic multiscale space-time lattice. The important question is whether this RG theory can be extended further to the fluctuating Navier-Stokes system? The obstacle for this development is the continuity of physical space and Galilean invariance. The other interesting direction of research is the study of spontaneous stochasticity in general, i.e. without a necessary connection to the Navier-Stokes system. To this end, the discrete-time symmetry-preserving models of this paper facilitate a mathematical formulation of the RG theory that is precise and convenient for numerical analysis.
@@ VIII.1 Implicit algorithm step
VIII.2 Anomalous scaling in discrete-time Sabra model
@@
VIII.3 Composition and sum of random maps

### v2

@@ VII Conclusion
The RG theory is formulated for discrete-time shell models. Such models are developed as approximations of the continuous-time Sabra model, converging to the latter as the approximation parameter $\varepsilon$ tends to zero. These discrete models are built on a dynamic space-time lattice and, from a numerical point of view, represent an adaptive finite-difference scheme. The key point is that these approximations exactly preserve scale symmetries and energy balance. It appears possible to extend such models to conserve other invariants, such as helicity; however, this does not seem to be important in the context of the present study.
The RG theory applies to class of models, in which the regularization (dissipation and noise) terms are scale invariant. The ideal limit corresponds to $N\to\infty$, where $N$ is a regularization (cutoff) parameter. For each $N$, the stochastic evolution is described by a flow kernel $\Phi^{(N)}$, which is a family of transition probabilities (Markov kernels) for all initial and boundary conditions. We derive the RG operator $\mathcal{R}$, which expresses the flow kernel $\Phi^{(N+1)}$ in terms of $\Phi^{(N)}$. This operator defines a dynamical system in the space of flow kernels and associates the ideal limit with an attractor of the RG dynamics. Here the dynamics (operator $\mathcal{R}$) depends on nonlinear couplings of the ideal system, but does not depend on the regularization (dissipation and noise). A specific regularization is introduced into the RG dynamics through the initial condition, i.e. through the initial flow kernel $\Phi^{(N_{0})}$.
@@
The RG approach to Eulerian spontaneous stochasticity was earlier developed for discrete models on a self-similar (fractal) lattice [28, 35]. Our current work demonstrates that the same approach works for systems with continuous time such as the Sabra shell model. This is achieved by using a dynamic multiscale space-time lattice. The important question is whether this RG theory can be extended further to the fluctuating Navier-Stokes system? The obstacle for this development is the continuity of physical space and Galilean invariance. The other interesting direction of research is the study of spontaneous stochasticity in general, i.e. without a necessary connection to the Navier-Stokes system. To this end, the discrete-time symmetry-preserving models of this paper facilitate a mathematical formulation of the RG theory that is precise and convenient for numerical analysis.
Finally, let us remark on another interesting role of noise by contrasting deterministic and stochastic regularizations. Consider deterministically regularized models, for example by setting $C=0$ in the LES shell model (17). One can then define a deterministic version of the RG operator acting on flow maps [36]. If such deterministic RG dynamics is chaotic in a classical sense, its invariant probability measure would yield a stochastic description in the limit $N\to\infty$. A natural question is whether this description coincides with the stochastic RG fixed point in Eq. (38). Although the answer may be positive, this analogy with chaos is not straightforward. First, the deterministic RG dynamics exhibits superexponential separation of trajectories [36], in contrast to the exponential growth characteristic of classical chaos. More importantly, probability evolution in the deterministic case is linear, as it is given by the pushforward of a deterministic map, whereas the stochastic RG operator in Theorem 1 is nonlinear, allowing, for example, periodic stochastic attractors [35]. These differences suggest a fundamental distinction between spontaneous stochasticity and deterministic chaos and point to an interesting direction for future research.
@@ VIII.1 Implicit algorithm step
VIII.2 Computer algorithm for the discrete-time model
Here we repeat the full algorithm described in Section III.1, reformulated in a way that is convenient for computer implementation. In the initialization step, one is given the cutoff shell $N$, the boundary state (9), the initial condition (10) and the terminal time $0<T<\infty$. Define the shell times $\tau_{n}=0$ for $n=1,\ldots,N$, with the respective variables $u_{n}=a_{n}$. The algorithm then proceeds by iterating the following steps:
  • (a)
    Choose $n$ as the smallest shell index for which $\tau_{n}$ attains the minimum among $\tau_{1},\ldots,\tau_{N}$.
  • (b)
    If $n=1$, then assign a new value to the variable $u_{1}$ according to Eq. (25).
  • (c)
    Choose time step $\Delta\tau_{n}$ from Eqs. (19) and (20) and update the time $\tau_{n}\mapsto\tau_{n}+\Delta\tau_{n}$.
  • (d)
    Compute the variables $u^{\prime}_{n}$ and $u^{\prime}_{n+1}$ according to Eq. (22); see also Appendix VIII.1. When $n=N$, the value $u^{\prime}_{n+1}$ is ignored; see Eq. (8).
  • (e)
    Assign a new value to the variable $u_{n}$ according to Eq. (24). Also, set $u_{n+1}=u^{\prime}_{n+1}$.
  • (f)
    The algorithm stops if $\tau_{n}=T$ for all $n=1,\ldots,N$. Otherwise, repeat the algorithm starting from step (a).
Note that the time steps are random, since they are expressed in terms of the turnover time (20), which depends on random state variables. This turnover time is bounded from below as
$T_{n}\geq\frac{1}{k_{n}}\big(\mathcal{E}+|b_{0}|^{2}+|b_{-1}|^{2}\big)^{-1/2},$ (58)
where $\mathcal{E}=\sum|u_{n}|^{2}$ is the energy. Because the time steps (19) are defined as an $\varepsilon$-fraction of $T_{n}$, it follows that the algorithm converges in a finite number of steps, provided that the energy does not blow up. From a physical point of view, it is natural to assume that a regularization is meaningful only if it does not lead to blowup; in this case, the algorithm converges in a finite—though random—number of steps.
Owing to the conservative form of the nonlinearity, the absence of blowup for the truncated fluctuating Sabra model (6) can be established by using the energy as a Lyapunov function; see, e.g., [37]. By the same reasoning, this global regularity is expected to extend to the corresponding discrete-time algorithm. On the other hand, the nonlinear form of the noise term in the LES shell model (17) may lead to blowup, and we observed numerical indications of such behavior when the noise coefficient $C$ is taken large compared to the dissipation parameter $D$. However, when $C$ is not large, no blowup behavior is observed numerically.
VIII.3 Anomalous scaling in discrete-time Sabra model
@@
VIII.4 Composition and sum of random maps
@@
Data availability. The data that support the findings of this article are openly available [10].